OpenStudy (anonymous):

jen and heather went to lunch. when the bill came jen discovered she had only $6 luckily heather had enough money to pay the other part, or 3/5 of the bill. how much did heather pay? how much was the bill?

OpenStudy (anonymous):

Let x be the bill x * 3/5 = 6 x = 6 * 5/3= 10 Heather paid 10-6 = $4 The bill was $10

OpenStudy (anonymous):

thanks